OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the y's to add out? 3x - y = 3 -2x + 2y = 6 2 3 none of the above 6 2. To solve this system using the addition method, you would need to multiply the first equation by what number in order for the x's to add out? 2x + 3y = 8 -6x + 4y = 11 6 3 none of the above -3 3. Solve the system using the addition method. 3x + 2y = 10 6x + 4y = 14 What is the solution to the system? (2, 2) Infinite solutions on the line No solution

OpenStudy (anonymous):

3. Solve the system using the addition method. 3x + 2y = 10 6x + 4y = 14 What is the solution to the system? (2, 2) Infinite solutions on the line No solution (1, 2) 4. Solve the system using the addition method. 3x + 4y = -10 5x - 2y = 18 What is the x-value in the solution? 2 none of the above -2 -4 5. Solve the system using the addition method. 5x - 2y = 3 2x - y = 0 What is the y-value in the solution? -3 6 3 none of the above

OpenStudy (anonymous):

You want someone to complete your entire assignment for you? That's not the way it works here unfortunately.

OpenStudy (anonymous):

no help

OpenStudy (anonymous):

helping you with no 3 only...

OpenStudy (anonymous):